The length of a rectangle is 7 feet more than its width. The length of a diagonal of the rectangle is 9 feet more than the width. By algebraic methods find the width of the rectangle.
Answer by ewatrrr(24785)


Hi,
Let w ,(w+7ft),(w+9ft)represent the width,length and diagonal respectively
Using the Pytahgorean Theorem
w^2 + (w+7)^2 = (w+9)^2
w^2 + w^2 + 14w + 49 = w^2 + 18w + 81
w^2 -4w - 32 = 0
factoring
(w-8)(w+4) = 0
(w+4) = 0
w = -4 Tossing out negative solution for length
(w-8) = 0
w = 8ft, the width of the rectangle (length=15ft, diagonal=17ft)
CHECKING our Answer
64 + 225 = 289